19. April 2016: Dienstag

LDAP/IdM:

  • Nightly Update: No new issues.
  • Accounts:
    • No errors in 29 new accounts.
    • Helped IAM with renames of multiple accounts

Backups:

  • Save Groups:
    • Five (5) clients that don’t backup because NetWorker messed up the aliases. Fixed and Re-run.
    • Failures: dpm1 and med23(O)
  • Clones: All still running.
  • Issues:
    • bujbod2 firmware update changed the info from the internal RAID1, sigh
    • Tape drive failure – new one arrived and was replaced.

SSL Certs – Mail

  • Investigated SSL certificate issue on the in-coming mail servers and why checktls.com was reporting failure.
  • Discovered that the server certs were replaced in March, but the new root and intermediate certificates from the CA were not installed at the same time.
  • Installed the current CA root and intermediate certs to match the new server certs on the in-coming mail servers

Tags:

Leave a Reply

You must be logged in to post a comment.